# Compressed data structures for Heegaard splittings

### v1

@@ 4. Algorithms on Heegaard diagrams
Referring to the discussion in the last paragraph of Section 2.1, one can find the fundamental group of $M$ from the Heegaard diagram data. Formally, given $I_{T}(\beta)$ delete, in linear time on the complexity of the SLPs, each reference of an edge $E-E_{T}(\alpha)$ and substitute each edge of $E_{T}(\alpha)$ by the symbol that represents its associated component, here seen as a generator $\pi_{1}(M)$. The total complexity is dictated by Theorem 3.3. ∎
@@
In a sense, we report Theorem 4.10 for the sake of completion; we are unaware of whether presentations of 3-manifold groups [3] where the relations are given by SLPs have any applications. Nevertheless, we note that, for example, the four Tietze transformations [30] are easily implemented in, at most, linear time on the SLPs complexity.
@@
The Smith normal form of the presentation matrix of $H_{1}(M)$ can be computed in time $\widetilde{O}(m^{2}g^{3}+m|T|)$, where $\widetilde{O}$ indicates some hidden polylogarithmic factors.
@@
We first compute the presentation matrix $K$ of the homology group, whose coefficients are algebraic intersection numbers between the diagram’s curves. Because we assume that each $a_{j}$ is a list of edges $E_{T}(a_{j})$, given an SLP $I_{T}(b_{i})$,
@@
which, by Lemma 2.2, can be computed in time $O(\|I_{T}(b_{i})\|\times\|E_{T}(a_{j})\|)$. Therefore, referring to Theorem 3.3 to compute $I_{T}(\beta)$, the total computational time of this procedure is $O\bigg{(}m|T|+\sum_{i=1}^{g}\sum_{j=1}^{g}\|I_{T}(b_{i})\|\|E_{T}(a_{j})\|\bigg{)}=O(m|T|+m|T|)$. Finally, for it to provide useful topological information, we need to reduce the matrix $K$ to its Smith normal form. This can be achieved in time $\widetilde{O}(g^{3}m^{2})$ by the Kannan-Bachem algorithm [53]. ∎
Theorem 4.13 (GET $\beta_{1}(M)$).
There is an algorithm to compute, in time $\widetilde{O}(mg^{3}+m|T|)$, the first Betti number $\beta_{1}$ of $M$.
Proof.
Given the Smith normal form outputted by GET $H_{1}$, first Betti number can be computed in time $O(g)$. However, if the Betti number is the only homological information required, rational Gaussian elimination can be used to obtain it in time $\widetilde{O}(g^{3}m)$ [4]. ∎
@@ 5. Examples and experiments
We here exemplify the usefulness of Heegaard diagrams by computing the homology group of manifolds in the so-called random Heegaard splitting model [3]. We say that the class of 3-manifolds is asymptotically $\mathcal{P}$, where $\mathcal{P}$ is a property of 3-manifolds (e.g., hyperbolic, Haken, etc.), if for all $g\geq 2$, letting $B(n)=\{\phi\in(\mathcal{L}\cup\mathcal{L}^{-1})^{*}||\phi|=n\}$,
$R=\lim_{n\to\infty}\frac{|\{\phi\in B(n)|(\Sigma_{g},\phi(\alpha))\text{ is }\mathcal{P}\}|}{|B(n)|}=1.$
Although it has been shown, for example, that manifolds are asymptotically hyperbolic [37, 38], rational homology spheres with torsion [14], and have particular probabilistic properties associated with some of their quantum invariants [15], other asymptotic properties, such as non-Hakeness and bounds on the hyperbolic volume [46], remain only conjectured.
We might as well consider an empirical version of the random splitting model, in which we independently draw $N$ words $({\phi}_{i})_{i=1}^{N}\overset{\mathrm{iid}}{\sim}\text{Uniform}(B(n))$ and consider the quantity
$\widehat{R}(n,N)=\frac{|\{\phi_{i}|(\Sigma_{g},\phi_{i}(\alpha))\text{ is }\mathcal{P}\}|}{N}.$
It is not hard to see that, as $n,N\to\infty$, $\widehat{R}(n,N)\to R$, meaning that the empirical model can be useful to explore some of the open questions regarding random splittings. Interestingly, however, to the best of our knowledge, there is a notable lack of studies which profit from this approach, with the few exceptions, such as [46], computing properties related to $\widehat{R}(n,N)$ by first triangulating Heegaard splittings with the Twister module [5] of SnapPy [11] and then using algorithms for triangulations to find relevant quantities. In some instances, such as for homology-dependent properties, the appeal to triangulations, nonetheless, not only adds computational cost but also seems to introduce some numerical instability, as we show in this section.
Here, we contrast Twister with a combination of the algorithm of Theorem 3.6 and GET $H_{1}$ (Theorem 4.12) to compute homology groups. Because the input of the experiments consists of Heegaard words, we note that there is no need to use proper normal coordinates or even triangulations of $\Sigma_{g}$; instead, we might compute the SLPs of the induced diagram with respect to general cellular embeddings, therefore avoiding the high cost of cyclic reductions.
@@ 5.1. Genus 1
Figure 6. Experiments with computing the homology group $H_{1}$ of Heegaard splittings over the torus. Frequencies and averages were computed for 30 repetitions.
Heegaard splittings over the torus are worth discussing not only because they yield a well-studied family of 3-manifolds, the lens spaces, but also because of the classical modular representation of $\operatorname{Mod}(\Sigma_{1})$ in $\text{SL}(\mathbb{Z}^{2})$ [19, 47], which allows for some correctness benchmarks to be traced. Explicitly, any Heegaard splitting $(T^{2},a,b)$ can be expressed through a vector in $(p,q)\in\mathbb{Z}^{2}$, where $p$ represents the number of signed intersections of $b$ and the usual longitude of the torus, $\ell$, and $q$ the signed intersections with the canonical meridian, $a$. Under this notation, Dehn twists about the meridian and longitude can be described by linear maps $(p,q)\mapsto(p,q-p)$ and $(p,q)\mapsto(p+q,q)$, respectively, and the homology group of $(T^{2},a,\phi(a))$ can be computed by $O(|\phi|)$-many $2\times 2$ matrix multiplications and a constant-time coordinate query, yielding an approach that we call the matrix method. In particular, direct recursion on the matrix method implies that homology group of the splitting of word $\phi_{n}=(\tau_{a}^{-1}\circ\tau_{\ell})^{n}$ equals $H_{1}^{n}=\mathbb{Z}/f_{2n}$, where $f_{n}$ is the $n$-th entry of the Fibonacci sequence starting at zero.
In Figure 6(a), the computed homology groups of the manifold $(T^{2},a,\phi_{n}(a))$ through Twister (with and without optimization, where by optimization the authors mean greedily folding of tetrahedra in the construction of a triangulation of $(\Sigma_{g},a,\phi(a))$, refer to their documentation [^2] Available at https://snappy.computop.org. for details) and the matrix method, implemented using numpy, are compared to the actual values $H_{1}^{n}$, computed from the Fibonacci sequences. These two methods are to be contrasted with ours. The experiments are repeated $30$ times for each value of $n$ to capture instabilities of SnapPy’s subroutines on the computation of $H_{1}$. The figure indicates the frequency which the homology groups computed are isomorphic to the expected $H_{1}^{n}$. Although all methods agree with the ground-truth homology $H_{1}^{n}$ for short sequences, when $n\geq 25$, Twister’s output is not always isomorphic to the expected group, while numpy deviates only after $n=47$. The lack of numerical robustness of these techniques might be traced to their reliance on rather unstable matrix manipulations, while our method, whose computed homology agrees with $H_{1}^{n}$ for all investigated values of $n$ (our method also agrees with $H_{1}^{n}$ for up to $n=500$, the maximum value we tested), uses only simple integer operations and therefore seems preferable for this problem.
In terms of computational time, our technique is slightly slower than Twister, as can be seen in Figure 6(b). Nevertheless, it is worth highlighting that, while the latter is partially implemented in C++, ours is fully in python. However, we often have significant gains over Twister in the general case, as shown in Figure 6(c), where we compare the average time to compute the first homology group of closed 3-manifolds defined by independently drawing random Heegaard words $({\psi}_{n})_{i=1}^{N}\overset{\mathrm{iid}}{\sim}\text{Uniform}(B(n))$ with $n$ now varying from $1$ to $500$. The speed-ups in computational time arise from simplifying repeated occurrences of Dehn twists in $\psi_{n}$ to power notation as in Theorem 3.6. While $\tau_{s}^{k}$ can be computed in time $\lg k$ (provided $k$ is given in binary) by our techniques, by triangulating Heegaard words with the current algorithms, one is forced to represent the manifold with at least $O(|\tau_{s}^{k}|)=O(k)$-many tetrahedra (before folding), meaning that the more repetitions in $\psi_{n}$, the greater is our gain with respect to Twister. This is stressed in Figure 6(d), where our computational time is compared to Twister’s (with optimization) for the family $(T^{2},a,\tau_{\ell}^{n}(a))$. There is some clear gain in our case: while the total complexity of Twister seems to be $O(n^{c})$ for some $c>1$, ours is only $O(n+\lg n)=O(n)$. What is more, if we allow inputs already in power-notation form, we can easily construct families of 3-manifolds for which the computation of their homology groups would require exponential time through Twister, but can be done in linear time in our framework.
@@
In fact, if we restrict the analysis to the very well-behaved layered triangulations [29], a triangulation of the lens space has a minimal of $\sum_{i}(a_{i})-1$ tetrahedra, where $(a_{0},\dots,a_{k})$ is the continued fraction expansion of $q/p$ (with $q$ and $p$ as defined for the matrix method). In particular, provided that we work solely with layered triangulations, we have a theoretical guarantee that $(T^{2},a,\psi_{2^{n}}(a))$ defines an infinite family of 3-manifolds that can be represented with Heegaard diagrams of complexity $n$, but with a triangulation of size at least $2^{n}$.
@@ 5.2. Genus 2
(a) The estimator $\overline{H_{1}^{n}}$ versus $n$.
@@
(b) Twister’s (with optimization) average computational time of $\overline{H_{1}^{n}}$ versus our method’s.
@@
Figure 7. Experiments with computing the estimator $\overline{H_{1}^{n}}$ for Heegaard splittings over the torus. Averages were computed for 30 repetitions.
For the case of Heegaard splittings over the double torus, no faithful representation of $\operatorname{Mod}(\Sigma_{2})$ is known, which means that we lack more effective methods to check the correct convergence of the homology computations. This time, however, we have that the random Heegaard splitting model asymptotically gives rational homology spheres with torsion [14], that is, of zero Betti number, but with $H_{1}$ finite. In particular, [46] shows that the random variable $\overline{H_{1}^{n}}=\frac{1}{n}\sum_{i=1}^{N}\log|H_{1}((\Sigma_{g},\psi_{n}^{i}(\alpha)))|$ where $(\psi_{n})_{i=1}^{N}\overset{\mathrm{iid}}{\sim}\text{Uniform}(B(n))$ follows a Central Limit Theorem
(5.2) $\frac{1}{\sqrt{N}}(\overline{H_{1}^{n}}-n\lambda)\to\mathcal{N}(0,\sigma^{2})$
for some positive constants $\sigma$ and $\lambda$, as $N\to\infty$ (restricted to cases where $\log|H_{1}((\Sigma_{g},\psi_{n}^{i}(\alpha)))|$ is finite). Equation (5.2) can be used as in Figure 7(a) to test the validity of our computations of the homology groups. In particular, we estimate the coefficient $\lambda$ as 0.13. For the genus $2$ case, Twister’s results and ours were this time identical for all experiments, but we were, once again, consistently faster in the average case, as indicated in Figure 7(b).
Remark 5.3.
For $g=2$, the homology groups of the manifolds $(\Sigma_{g},\alpha,\psi_{n}(\alpha))$ estimated by Twister and our methods agreed up to $n=1000$, which includes groups of order significantly superior to the orders that both methods started to deviate in the $g=1$ case. This implies that the inconsistencies described in Section 5.1 are likely not due to some difference in the maximum integer storable in the implementation of Twister and ours, as one might be inclined to suppose. Moreover, to guarantee that the errors did not arise from the subroutines where we computed Smith normal forms—which are used in the case of $g=2$, but unnecessary for lens spaces—we experimented with different implementations of the Kannan-Bachem algorithm, namely sympy’s [41] and SNF’s [40], all which were consistent with SnapPy’s implementation. The lower robustness of Twister in computing the homology of lens spaces and its behavior for $g>2$ remains therefore to be investigated.

### v2

@@ 4. Algorithms on Heegaard diagrams
Given $I_{T}(\beta)=\{I_{T}(b)\}_{b\in\beta}$ delete, in linear time on the complexity of the SLPs, all references to the edges $E-E_{T}(\alpha)$ and substitute each edge of $E_{T}(\alpha)$ by the symbol that represents its associated component $a\in\alpha$, here seen as a generator $\pi_{1}(M)$. This gives $g$ distinct SLPs $r_{i}=[x_{1}^{i}=\text{EXPR}_{1}^{i},\dots,x_{\ell_{i}}^{i}=\text{EXPR}_{\ell_{i}}^{i}]$ whose complexities add up to $m$ and describe the relations of the presentation of $\pi_{1}(M)$. A presentation of $\pi_{1}(M)$ can then be defined with generators $a_{1},\dots,a_{g},x_{1}^{1},\dots,x_{\ell_{i}-1}^{1},\dots,x_{\ell_{g}-1}^{g}$ and relators $(x_{j}^{i})^{-1}\cdot\text{EXPR}_{j}^{i}$ (this corresponds to repeated applications of a Tietze move [johnson1997presentations] to the presentation of $\pi_{1}(M)$ computed using the method described at the end of Section 2.1). The output is a balanced presentation of $\pi_{1}(M)$ with $\|I_{T}(\beta)\|$-many generators and relators. The total time is dictated by Theorem 3.2. ∎
@@
Because the outputted presentation of $\pi_{1}(M)$ in Theorem 4.10 has linear size on the input, most complexity results about uncompressed presentations of 3-manifolds (such as triangulations) that appeal to the fundamental groups still hold in our compressed data structure. This is the case, for example, of Zentner’s proof that sphere recognition lies in co-NP [Zentner_2018] (assuming the generalized Riemann hypothesis). Similarly, the usual bounds on the volume [cooper1999volume] and diameter [white2001diameter] of hyperbolic 3-manifolds can be naturally translated to our notion of complexity.
@@
The Smith normal form of the presentation matrix of $H_{1}(M)$ can be found in time $\widetilde{O}(m^{2}g^{3}+m|T|)$, where $\widetilde{O}$ indicates some hidden polylogarithmic factors.
@@
We first compute the presentation matrix $K$ of the homology group whose coefficients are algebraic intersection numbers between the diagram’s curves. Because we assume that each $a_{j}$ is a list of edges $E_{T}(a_{j})$, given an SLP $I_{T}(b_{i})$,
@@
which, by Lemma 2.3, can be computed in time $O(\|I_{T}(b_{i})\|\times\|E_{T}(a_{j})\|)$. Therefore, referring to Theorem 3.2 to compute $I_{T}(\beta)$, the total computational time of this procedure is $O\bigg(m|T|+\sum_{i=1}^{g}\sum_{j=1}^{g}\|I_{T}(b_{i})\|\|E_{T}(a_{j})\|\bigg)=O(m|T|+m|T|)$. Finally, we reduce the matrix $K$ to its Smith normal form in time $\widetilde{O}(g^{3}m^{2})$ by the Kannan-Bachem algorithm [sergeraert2024kannanbachemalgorithm]. ∎
@@ 5. Examples and experiments
Given a Heegaard word presentation $(\Sigma_{g},\phi)$ of a closed 3-manifold $M$, there exist algorithms to compute, in linear time on the length of $\phi$, a triangulation of $M$ with as many tetrahedra [alagic2011quantum, alagic2014quantum, brinkmann2001computing, koenig2010quantum]. This implies, for example, that the first homology group of $M$, $H_{1}(M)$, can be computed in polynomial-time on $|\phi|$. These algorithms should be contrasted with Theorem 3.5, which says that SLPs of the intersection sequence $I_{T}(\beta=\phi(\alpha))$ can be computed in linear time on the complexity of $\phi$ measured in power-notation form. Because the fundamental group of $M$ depends only on $I_{T}(\beta)$, algebraic information on splittings with significantly complicated gluing maps can be retrieved by Theorems 4.12 and 4.10.
Here, we explore the gains from the compressed data structure by comparing an early implementation of our methods with SnapPy’s [SnapPy] Twister [Twister] module, commonly used to compute triangulations from Heegaard words. Since experiments are limited to Heegard words as input, there is no need to use proper normal coordinates or even triangulations of $\Sigma_{g}$; instead, we may compute the SLPs of the induced diagram with respect to general cellular embeddings, therefore avoiding the high cost of cyclic reductions. This guarantees that, for the problems analyzed, our theoretical complexity is at least equal to Twister’s. Even without these simplifying assumptions, we are still exponentially faster in the worst cases, as we show in the following section about lens spaces. Similar computational gains for genus 2 splittings are shown in Section 5.2, where some statistical properties of the homology groups of random splittings [rivin2014statistics] are explored.
@@ 5.1. Genus 1
Figure 9. Experiments with computing the homology group $H_{1}$ of Heegaard splittings over the torus. Averages were computed for 30 repetitions.
Heegaard splittings over the genus 1 surface—i.e., the torus—form the well-known family of lens spaces. Every lens space is uniquely given by a pair of integers $(p,q)$, which, for a Heegaard splitting $(\Sigma_{g},a,b)$, represents the algebraic intersection numbers of $b$ with the usual longitude and meridian of the torus, respectively (there is some ambiguity in the values of $p,q$, see [saveliev2011lectures]). A Dehn twist about the meridian $m$ of the torus maps $(p,q)$ to $(p+q,q)$, while about the longitude $\ell$ maps to (p,qp)(p,q-p). This gives a representation of $\operatorname{Mod}(\Sigma_{1})$ in $\mathbb{Z}^{2}$ from which the homology of the Heegaard splitting $(\Sigma_{g},\phi)$ can be exactly computed by $|\phi|$ matrix multiplications and one query on the final value of $q$. We used this approach to benchmark the accuracy of our method, which returned the expected $H_{1}$ for input words $\phi$ of length up to 1000, the maximum size we tested.
Figure 9 compares the average computational time needed to find $H_{1}(\Sigma_{1},\phi)$ taken by our method and by Twister for four different families of $\phi$ indexed by an integer $n$: (a) words of form $(\tau_{m}\tau_{\ell}^{-1})^{n}$; (b) uniformly sampled words of length $n$; (c) words of form $\tau_{\ell}^{n}$; and (d) words of form $\tau_{\ell}^{n}$, but this time, given in power-notation. Averages were taken over 30 repetitions. In all cases, we compared our times with Twister’s, with and without optimization (where, by optimization, the authors mean the greedy folding of tetrahedra; refer to their documentation for details [^2] Available at https://snappy.computop.org.). Although optimization drastically reduces computational time, it seems to increase numerical error. In every instance, except for (d), we first transformed the words into a power-notation form before using them to compute $I_{T}(b)$; in total, this can be done in linear time on $|\phi|$.
@@
A direct application of the representation of $\operatorname{Mod}(\Sigma_{1})$ implies that, for the family of case (a), $H_{1}^{n}=\mathbb{Z}/f_{2n}$ where $f_{n}$ is the $n$-th Fibonacci number (starting at 0). We used this relation to benchmark the accuracy of the methods for different values of $n$. While our technique always agrees with the expected value for up to $n=500$, Twister (with optimization both on and off) starts to deviate for values of $n\geq 25$. Interestingly, we did not observe divergences for higher genii splittings.
Case (a) is where our advantages compared to Twister are the worst. This is because the words $(\tau_{m}\tau_{\ell}^{-1})^{n}$ do not admit a non-trivial power-notation form, which is mainly responsible for our gains in the other cases. Nevertheless, it is worth highlighting that, while Twister is partially implemented in C++, our code is fully python. Case (b) represents the average case of computing the homology of lens spaces given a gluing map $\phi$. We notice that we are now consistently faster for all values of $n$ tested. This happens exactly as we explore redundancies by first putting the word in power-notation form, with significant gains every time the same Dehn twist occurs in sequence (since $\operatorname{Mod}(\Sigma_{g})$ has two generators, this will be the case quite often). The gains due to repetitions are further highlighted in (c): while Twister is bound to take $O(n^{c})$, for some $c>1$, to compute the homologies, our approach takes time $O(n+\lg n)$, where the linear dependence comes solely from compressing the inputs to power-notation. In fact, if we assume the words already in power-notation, we are exponentially faster than Twister, as shown in case (d).
We can explore our structure’s usage of power-notation to compute invariants of families of 3-manifolds that would necessarily require exponentially many tetrahedra to be represented. The minimum number of tetrahedra in a layered triangulation of a lens space $L(p,q)$ is $\sum_{i}(a_{i})-1$, where $(a_{0},\dots,a_{k})$ is the continued fraction expansion of $q/p$ [jaco2006layered]. Moreover, [jaco2009minimal, jaco2011coverings] discovered families of lens spaces—including the spaces defined by $p=2n,q=1$ and $p=4n,q=2n-1,n\geq 2$—for which layered triangulations are minimal among all triangulations. So, for example, while the family $L(2^{n},1)$ necessarily requires exponentially many tetrahedra to be represented, it can be encoded in linear space in our data structure. It seems natural to expect that there are many other families of manifolds, with diagrams exponentially smaller than the minimum number of tetrahedra, for which our techniques can provide real computational gains compared to any existing method.
@@ 5.2. Genus 2
For the case of Heegaard splittings over the double torus, there is no easy version of the representation trick to compute the homology group, which means that we lack effective accuracy benchmarks. However, this time, we can consider a random Heegaard splitting model [aschenbrenner20153]. Formally, suppose we uniformly draw words from the radius-$n$ ball $B(n)=\{\phi\in\operatorname{Mod}(\Sigma_{g})||\phi|=n\}$ for a fixed genus $g\geq 2$. We say that the class of 3-manifolds is asymptotically $\mathcal{P}$, where $\mathcal{P}$ is a property of 3-manifolds (e.g., hyperbolic, Haken, etc.), if, for all $g\geq 2$, the probability that a manifold with splitting $\phi\sim\text{Uniform}(B(n))$ has the property $\mathcal{P}$ tends to 1 as $n\rightarrow\infty$. While it has been shown, for example, that manifolds are asymptotically hyperbolic [lubotzky2016random, https://doi.org/10.1112/jtopol/jtq031], rational homology spheres with torsion [dunfield2006finite], and have particular features associated with some of their quantum invariants [dunfield2011quantum], other asymptotic properties, such as non-Hakeness and bounds on the hyperbolic volume [rivin2014statistics], remain only conjectured.
(a) The estimator $\overline{\log H_{1}^{n}}$ versus $n$, computed using our method.
@@
(b) Twister’s (with optimization) average computational time of $\overline{\log H_{1}^{n}}$ versus our method’s.
@@
Figure 10. Experiments with computing the estimator $\overline{\log H_{1}^{n}}$ for Heegaard splittings over the double torus. Averages were taken over 30 repetitions.
Interestingly, although the random model is naturally suitable for empirical study, to our knowledge, there is a notable lack of works that profit from this approach, with one prominent exception [rivin2014statistics]. In this paper, Rivin shows that the variable $\overline{\log H_{1}^{n}}=\frac{1}{n}\sum_{i=1}^{N}\log|H_{1}((\Sigma_{g},\phi_{n}^{i}(\alpha)))|$ where $\{\phi_{n}^{i}\}_{i=1}^{N}\overset{\mathrm{iid}}{\sim}\text{Uniform}(B(n))$ follows a Central Limit Theorem
(5.2) $\frac{1}{\sqrt{N}}(\overline{\log H_{1}^{n}}-n\lambda)\rightarrow\mathcal{N}(0,\sigma^{2})$
for some positive constants $\sigma$ and $\lambda$, as $N\rightarrow\infty$ (restricted to cases where $\log|H_{1}((\Sigma_{g},\phi_{n}^{i}(\alpha)))|$ is finite). In particular, he uses Twister to verify this relation for the case of $g=2$. In Figure 10(a), we reproduce the analysis and estimate $\lambda$ as 0.13. This time, we are about 2.5 times faster than Twister in the average case (10(b)). That both methods converge to the expected regression line is a strong indication of the correctness of the implementations.
(a) The estimator $\overline{\log H_{1}^{n}}$ versus $n$ where the words sampled are $\phi_{n,r}$.
(b) The estimated coefficients $\lambda$ as a function of $r$.
(c) The estimator $\overline{\log H_{1}^{n}}$ versus $n$ where the words sampled are $\widetilde{\phi}_{n}$.
(d) The estimated geometric mean of $H_{1}^{n}$ for different $n$.
Figure 11. Experiments with computing the estimator $\overline{H_{1}^{n}}$ for Heegaard splittings over the double torus with different sampling strategies. Averages were taken over 30 repetitions.
Due to the efficiency of our technique for input in power notation, we are naturally inclined to consider modifications to the sampling strategy used to compute $\overline{\log H_{1}^{n}}$ by drawing words from different distributions. For example, Figure 11(a) shows the estimator $\overline{\log H_{1}^{n}}$ as a function of $n$, where we now draw words $\{\phi_{n,r}^{i}\}_{i=1}^{N}$ uniformly from $\mathcal{L}\times\llbracket-r,r\rrbracket$ for different choices of $r$. The distribution of $\phi_{n,1}$ is identical to $\phi_{n}$, described in the previous paragraph, and consequently, follows the linear relation described by Rivin. However, it is noticeable that the relationship between $n$ and $\overline{\log H_{1}^{n}}$ is still linear, even for different values of $r$, although $\lambda_{r}$ grows as $\log(r)$ (Figure 11(b)). This logarithmic relationship should come as no surprise, given the direct dependency of $H_{1}$ on the minimal number of intersections between $\alpha$ and $\beta$ curves up to isotopies.
A more interesting modification to the sampling technique, this time denoted by $\{\widetilde{\phi}_{n}^{i}\}$, yields Figure 11(c). For $\{\widetilde{\phi}_{n}^{i}\}$, we fix a value of $n$ and sample $(\tau_{s_{j}},k_{j})$ uniformly from $\mathcal{L}\times\llbracket-n_{j},n_{j}\rrbracket$, with the constraint that, if the word sampled equals $\tau_{s_{\ell}}^{k_{\ell}}\circ\dots\circ\tau_{s_{1}}^{k_{1}}$, then $\sum_{j}|k_{j}|=n$. This distribution aims to mimic the random model, giving higher weights to words with more frequent repetitions. As indicated by the figure, $n$ and $\overline{\log H_{1}^{n}}$ seem to follow a logarithmic relation. The estimated model is shown in red; its $R^{2}$ was found to be $0.86$. Due to the logarithmic form, one is inclined to conjecture that there is a power-law relation between the geometric mean of $H_{1}^{n}$ and $n$ when sampled according to $\{\widetilde{\phi}_{n}^{i}\}$; nonetheless, as indicated by Figure 11(d), even if such a relation exists, the distribution has heavy tails. We leave for future work a theoretical justification for these relations.
6. Open questions towards a census of 3-manifolds
The strategy for generating splittings used in Section 5 suggests a natural construction of a census of closed 3-manifolds containing elements that would otherwise be inaccessible to triangulation-based approaches. Being more explicit about the “recipe” presented in the introduction, one could
  1. (1)
    Fix a genus $g$ and generate Heegaard words in power-notation form up to a maximum length, which are then used to compute diagrams;
  2. (2)
    Greedly apply combinatorial moves (such as GET EFFICIENT POSITION, DO A DISK SLIDE, DO DESTABILIZATION, and DETECT REDUCTION) to simplify as much as possible the diagrams and detect some trivial, previously identified instances, such as stabilizations;
  3. (3)
    Use 3-manifold invariants to detect redundancies in the enumeration of spaces.
A few points about these strategies deserve mention. Instead of sampling Heegaard words as in item 1, a more natural (and less studied) model would come from directly sampling the diagrams as normal coordinates, without regard for an input word. Although normal coordinates respect some well-defined combinatorial constraints, efficiently restricting the multicurves to systems in the surface (without, for example, appealing to Theorem 4.1) does not seem to be an easy task. Similar problems appear when sampling 3-manifolds from random triangulations, for which some asymptotic results are already known [petri2022model].
Now for item 3, we note that the very efficient computations of (presentations of) the fundamental and homology groups in our data structure represent some advances, but the difficulty of the isomorphism problem for the first [aschenbrenner20153] and the lack of strong distinguishability of the second make them not ideal in most cases. Alternatively, the authors of [costantino2020kuperberg] suggest a presentation of the much distinguishable Kuperberg [kuperberg1991involutory] and Turaev-Viro [turaev2010quantum] invariants computed directly from Heegaard diagrams of 3-manifolds, but an algorithm on our data structure and its complexity for the cases of interest still need to be determined.
Finally, in addition to the exploration of exponentially more spaces, a Heegaard splitting census can probably be used to address the well-reported difficulty of generating 3-manifolds with particular topological structures—such as hyperbolicity—observed in triangulation-based methods [matveev2007algorithmic]. This is because much of the structure of a 3-manifold is encoded by the Hempel distance of its splittings, defined as a graph distance in the infinite complex of meridians of the involved handlebodies, loosely related to the number of intersections between $\alpha$ and $\beta$ [hempel20013]. For example, manifolds that have splittings with a Hempel distance of at least 1 are irreducible [kobayashi1988casson], distance of at least 3 are hyperbolic [hempel20013], and distance of at least $k$ cannot embed incompressible orientable surfaces of the genus at most $2k$ [hartshorn2002heegaard]. Bounds relating the hyperbolic volume [hamenstadt2022stability, kobayashi2013hyperbolicvolumeheegaarddistance] and the Heegaard genus [scharlemann2006alternate] with the Hempel distance are also known. There exists a well-known algorithm, due to Masur and Schleimer [masur2013geometry], to approximate the Hempel distance from a Heegaard diagram, which can probably be translated into our structure, but whose complexity is not clear. Alternatively, the authors of [ennes2025hardnesscomputationquantuminvariants] give a polynomial-time algorithm that uses (uncompressed) intersection sequences of Heegaard splittings to increase the Hempel distance while keeping some quantum invariant fixed throughout the process, making it therefore possible to automatically generate sets of hyperbolic manifolds that are guaranteed not to be homeomorphic.